Show detailed solution k = 9 Times 10^9 Nm^2/C^2,
Answer #1

a)

agram below, the Clai ge a 23 3 13 5.0 cm q1 8.0 cm 2

Sorry for the design. I know I should refine a bit, but the time is much more important (for me) than such a detail.

b)

F13=kq1q3/r132

F13~3.54*10-9 N

c)

F23=kq2q3/r232

F23=16.2*10-9 N

d)

Ox: Rx=F13x=F13cos(alpha)~3.002*10-9 N

Oy: Ry=F23-F13y=F23-F13sin(alpha)~14.324*10-9 N

With:

cos(alpha)=0.848

sin(alpha)=0.53

R=\sqrt{R_{x}^{2}+R_{y}^{2}}

R~14.635*10-9 N.

e)

The angle between R and Ox axis is: tan(beta)=Ry/Rx ---> beta~-78.16 deg.

If counterclockwise from Ox: ~101.84 deg.

Please, check the numerical calculation.